OpenStudy (anonymous):

A particle moves with a constant acceleration of (0.1i-0.2j)ms^-2 it is initially at the origin where it has velocity (-i+3j)ms^-1. i and j are east and north Find the speed of the particle when it is traveling south east

OpenStudy (michele_laino):

using this reference frame:  the speed of the particle is given by the subsequent equations: \[\Large \left\{ \begin{gathered} {v_x}\left( t \right) = 0.1t - 1 \hfill \\ {v_y}\left( t \right) = - 0.2t + 3 \hfill \\ \end{gathered} \right.\]

OpenStudy (michele_laino):



OpenStudy (michele_laino):

for example, at t=10 seconds, we have: \[\Large \left\{ \begin{gathered} {v_x}\left( {10} \right) = 0.1 \times 10 - 1 = 0 \hfill \\ {v_y}\left( {10} \right) = - 0.2 \times 10 + 3 = 1 \hfill \\ \end{gathered} \right.\]

OpenStudy (anonymous):

Thanks, this is using a simplified version of the equation v=u+at?

OpenStudy (anonymous):

and would it be correct to say that when it is travelling south east i is equal to -j ?

OpenStudy (michele_laino):

yes! I have used that formula, and south east means that our particle has the speed oriented as -j and i

OpenStudy (anonymous):

Thanks i got the answer, this helped a lot.

OpenStudy (anonymous):

Aww isn't this adorable

OpenStudy (anonymous):

nice wolf

OpenStudy (anonymous):

Say Robdogg, did you ever clear up them herpes?